OpenStudy (anonymous):

A computer company gives a discount of 10 % on a new piece of software. If it makes a profit of 7%, what percent does the marked price exceed the cost price?

OpenStudy (anonymous):

let marked price is y and cost price=x. y*90/100 =x*107/100; y/x=107/90 (y-x)/x=17/90; y-x 170 ----*100% =------% =18.89% x 9